1) A ball moves with velocity v=(2, 3.4, 6) m/s (meter per second). At time t = 0, it is located at position p=(0, 5, 2).
  
   a) Find the equation of its path in parametric form.

  
  
   b) At time t = 10 second, what is the ball's position?
Answer #1

a)

r(t) = r_0 + tv = (0,5,2) + t(2,3.4,6) = (0+2t,5+3.4t,2+6t) = (2t,5+3.4t,2+6t)

b)

When the time is equal to 10 seconds, the position of the ball will be given by r(10), which will be

r(10) = r_0 + 10v = (0,5,2) + 10(2,3.4,6) = (0+2(10),5+3.4(10),2+6(10)) = (20,39,62)

Note - Post any doubts/queries in comments section.